OpenStudy (anonymous):

y = −2x2 + 8x − 12 Find the vertex :)

OpenStudy (anonymous):

Please help me :)

OpenStudy (anonymous):

Complete the square to find the transformation. -2(x-2)^2-4 the y value is -4. so the vertex is at (2,-4) OR Use the vertex formula: -b/2a a=-2 b=8 x=-8/(2)-2 x=2 Run 2 through the original equation: y=-(2)^2+8(2)-12 y=4

OpenStudy (anonymous):

Thanks ;)